OpenStudy (calculusxy):

MEDAL! Abe can mow the lawn in 15 min. Ben can mow the lawn in 20 min. Carl can mow the lawn in 30 min. They work together for 5 min. What part of the lawn was mowed? @phi

OpenStudy (phi):

I think of these as rate* time= distance, except distance is really Jobs for example Abe works at a rate of 1 lawn per 15 min or 1/15 now figure out the rates for all 3, add them up and multiply by time (5 min) to get the fraction of a job they do.

OpenStudy (ranga):

In five minutes, Abe mows 5/15th of the lawn Ben mows 5/20th of the lawn Carl mows 5/30th of the lawn Together they mow: 5/15 + 5/20 + 5/30 of the lawn. Add them up. That is the part of the lawn they would have mowed in 5 minutes.

OpenStudy (calculusxy):

@ranga Can I also simplify then add?

OpenStudy (ranga):

Yes. That is what I would do.

OpenStudy (calculusxy):

1/3+1/4+1/6=4/12+3/12+2/12=9/12=3/4 Am I right?

OpenStudy (ranga):

Yes.